A person takes a trip, driving with a constant speed of 99.0km/h except for a 22.0-min rest stop. The person's average speed is 73.1km/h.

(a) How much time is spent on the trip?

(b) How far does the person travel?
Answer #1

Driving speed u = 99km/hr

Average speed ,Vav- 73.1km/hr

22 min stop means 22min/600.367hr

---------

Time spent on trip =Time spent on driving +time spent on resting

T=t_{d}+t_{r}

Speed =Distance/time

Distance =Speed *time

Total distance driven =Driving speed* time spent on driving

D=v*t_{d}

D = 99km / hr * td

-----------

Average Speed =Total Distance/Total time

Unu-

73.1km/hr-(99km/hr *ta)/(ta tr)

73.1 (99 * ta)/(4 + 0.367)

73.1 (4 + 0.367) = 99 * td

73.1 * td 73.10.36799 td

73.1 * 0.367 = (99-731) * td

73.1*0.367/25.9=t_{d}

{color{Red} t_{d}=1.036hr}

T=t_{d}+t_{r}=1.036hr+0.367hr

(a)ANSWER: {color{Red} T=1.403hr}

=====================

(b) How far does the person travel?

D=v*t_{d}

D=99km/hr*1.036hr

ANSWER: {color{Red} D=102.564km}

===========================